### v1

# Which chromium-sulfur compounds exist as 2D material?

Abstract
Two-dimensional (2D) chromium-sulfides are synthesized by molecular beam epitaxy using graphene as a substrate. Structure characterization by employing scanning tunneling microscopy and low energy electron diffraction indicates that there are two 2D phases, Cr2S3-2D and CrS4-2D, which have not been reported before. CrS4-2D is related to bulk Cr5S6, but thinner than a bulk unit cell. For Cr2S3-2D, an even thinner material, no bulk counterpart exists. Both 2D materials are found to be structurally stable under ambient conditions and exhibit interesting electronic properties. Extensive first-principles calculations provide further insight into the electronic structure of these systems and indicate that they should be magnetic. Surprisingly, CrS2 recently reported from density functional theory calculations and experiments could not be fabricated.
1 Introduction
Research on chromium-chalcogene compound 2D materials has recently flourished, with significant advancements both theoretically in the prediction of phases and properties as well as experimentally in the preparation and investigation of new phases 1, 2, 3. One element of excitement, shared with chromium halides 4, 5, 6, is the interesting magnetic properties of these new 2D materials 7, 8, 9, 10, 11, 12, 13, 14, 15. Focusing on the case of chromium-sulfur 2D materials 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41 it seems not yet to be settled which phases exist as 2D materials. Neither were all compounds predicted from DFT calculations experimentally fabricated nor is the experimental research comprehensive enough to make additional discoveries unlikely. This situation is one of the motivations for the present study.
While the structure and composition of 2D materials created by exfoliation are known from their bulk crystal, the situation is far more intricate for 2D materials grown by molecular beam epitaxy (MBE) or chemical vapor deposition (CVD). These 2D materials may possess no bulk counterpart; thus, their structure and composition are not known a priori. The thinness and lack of signal make standard techniques like X-ray diffraction or Raman spectroscopy difficult to apply. The unknown arrangement of planes of different species, depth-dependent attenuation, and vacancy-containing planes impair even composition analysis methods like X-ray photoelectron spectroscopy (XPS). Frequently, density functional theory (DFT) calculations are used as a guideline for interpreting experimental compounds, although stability in DFT calculations does neither guarantee a pathway for creation nor the existence of a specific phase. All of this is relevant to 2D materials based on chromium and sulfur.
The bulk chromium sulfur phase diagram is complex 42, 43, 44, 45, but no Cr-rich phases have been documented. Unquestioned S-rich phases are limited in stoichiometry by Cr2S3, except of the more S-rich Cr5S8 phase which can be prepared under high pressure. Noteworthy, no compound with van der Waals gaps has been reported as bulk phase.
Bulk Cr2S3 can be considered to be composed of hypothetical CrS2 layers with octahedral (T) Cr coordination, where the interlayer gaps are filled by 1/3 of the possible sites establishing chemical bonds between the CrS2 layers. Depending on the precise conditions of preparation and composition, the distribution of interlayer Cr atoms may be disordered or give rise to trigonal or rhombohedral stacking [see Figure 1(a)] 43, 46, 47. For the latter two, the arrangement of the interlayer Cr repeats every two or three CrS2 layers, respectively. Complex magnetically ordered structures are reported, which critically depend on the arrangement of Cr interlayer atoms and the exact composition 46, 48, 49, 47, 50. The magnetic moments are in-plane, normal to the c-axis 48, 47. Rhombohedral Cr2S3 is a semiconductor, paramagnetic above and ferrimagnetic below 120 K with a maximum magnetization around 90 K 49, 46, 47, 50. Also negative magnetostriction and photoconductivity were reported for rhombohedral Cr2S3 50, 51.
Up to now, only the rhombohedral variant of Cr2S3 has been synthesized as 2D material 16, 17, 18, 19, 20, 21, 22, 23, 24, using either chemical vapor deposition or sulfurization of a metallic film. Single unit cell thickness has been reported 16, 17, 19, 22, 24, with the unit cell consisting of four CrS2 layers with three planes of 1/3 Cr sites occupied in the interlayer gaps. However, in a transmission electron microscopy study the arrangement of Cr interlayer atoms in as-grown "rhombohedral" CVD films was found to be random 22. Similar to bulk counterpart, semiconducting behavior, magnetization with a T K, negative magnetostriction, and good photosensitivity were found 17, 20, 19.
While no bulk compound of composition CrS2 has been reported, DFT calculations proposed the existence of CrS2 as a typical transition metal dichalcogenide with the S atomic planes sandwiching the Cr atomic plane [see Figure 1(b)] 37, 36. H- (trigonal prismatic Cr coordination), T-, and T’- phases were identified as possible variants, with ambiguity on the most stable phase being either the H- 37, 31 or the T-phase 36. Recently, also two different orthorhombic CrS2 phases were identified in calculations 33, 52. DFT based research for CrS2 materials is flourishing due to exciting predicted properties ranging from ferromagnetism 36, 33, 31 with Curie temperatures up to 1000 K, over a direct band gap with valley polarization 37 to excellent performance for the hydrogen evolution reaction 35, to name a few.
Despite the absence of the bulk compound, the synthesis of CrS2 as 2D material has been communicated25, 30, 27, 26, 28, in the T-phase 25, 27, 26, 28, 40, 39, but also in the H- and T’-phases 25. CrS2 islands tend to grow thicker than single unit cell, but in two cases minimal heights in the range of nm were measured 25, 28. Transport in CrS2 field effect devices was characterized as p-type 25, n-type 26, or metallic 27, 28. Weak ferromagnetic behavior up to room temperature 27 and use as a floating gate in photoelectric non-volatile memory 28 were described.
In bulk Cr5S6, 2/3 of the possible Cr sites in the interlayer gaps are filled, i.e., it is a material in the NiAs structure with 1/3 of Cr vacancies in every second Cr layer [see Figure 1(c)] 43. Cr5S6 is antiferromagnetic below 168 K and ferrimagnetic up to 303 K 48. Up to now, to the authors’ best knowledge, preparation of Cr5S6 as a 2D material has not yet been reported.
Lastly, there is a yet unconfirmed DFT prediction for a new CrxSy 2D material, namely Cr2S3-2D [see Figure 1(d)] 40, 39. To avoid confusion with bulk materials with the same composition but different structure, "-2D" is attached to the stoichiometric formulas in the 2D case from hereon. Cr2S3-2D is unrelated to bulk Cr2S3 and consists of five planes of atoms in the sequence S-Cr-S-Cr-S in the NiAs structure, which can be related to bulk CrS, although it lacks the monoclinic distortion observed there 43. Based on the DFT calculations, Cr2S3-2D is proposed to be a magnetic semiconductor with a transition temperature above room temperature (405 K) 53. The key difference of Cr2S3-2D compared to other 2D magnetic semiconductors, such as CrI3, is the presence of both intralayer and interlayer exchange coupling 40. This characteristic of Cr2S3-2D may account for its high transition temperature.
Given the diversity of materials and predictions, it remains worthwhile to carefully characterize and investigate CrxSy-2D phases. In this paper, we use molecular beam epitaxy (MBE) to prepare CrxSy-2D on graphene and subsequently characterize the materials in situ by low energy electron diffraction (LEED), scanning tunneling microscopy (STM) and spectroscopy (STS). Two phases can be prepared phase pure and are identified as Cr2S3-2D [Figure 1(d)] and CrS4-2D [see Figure 1(e)]. While Cr2S3-2D has already been predicted to exist as mentioned above, CrS4-2D lacks such a DFT based forecast. It can be considered the minimum thickness variant of bulk Cr5S6, i.e., it consists of two CrS2 layers with additional Cr filling 2/3 of the sites in the interlayer gap. Moreover, based on our experiments, we tentatively assert that the existence of CrS2 as 2D-material is questionable, consistent with the non-existence of this material in bulk.
2 Results and discussion
2.1 Structure from experiment
CrxSy-2D compounds are grown on single crystal graphene (Gr) prepared on Ir(110). For the LEED pattern in Figure 2(a) after deposition of Cr in S vapor at room temperature, the sample was post-annealed in S vapor at 750 K (see Methods for details). Besides the Gr/Ir(110) substrate reflections 54, a diffraction ring is also visible. It has a sixfold symmetric intensity modulation with the highest intensity aligned with the first-order Gr reflections. We assign this diffraction ring with radius to the newly formed CrxSy-2D compound. Based on the sixfold symmetry of the diffraction ring, a threefold or sixfold in-plane symmetry of the new compound can be assumed, resulting in an in-plane lattice constant of nm.
The typical topography of the sample is presented through the STM image of Figure 2(b). Gr/Ir(110) is covered with an area fraction of about by coalesced CrxSy-2D islands with wavy edges and typical linear dimensions of 50 nm. Faint dark lines are visible within the islands, which we interpret as grain boundaries formed during annealing due to the coalescence of almost randomly oriented smaller islands. They are best visible in the contrast-enhanced box in Figure 2(b). There, also the larger nm Gr/Ir(110) moiré periodicity can be recognized which is imprinted from the substrate onto the island. The atomically-resolved STM image of CrxSy-2D in the inset displays a surface lattice of periodicity of nm consistent with determined from LEED. As shown by the height profile along the white line in Figure 2(b), the islands display a uniform apparent height of nm at a sample bias V. Irrespective of the precise preparation conditions, we never observed islands of lower height. As justified below, we interpret this CrxSy-2D compound as Cr2S3-2D of which the side view ball model is shown in Figure 1(d).
Annealing to successively higher temperatures using the same sulfur pressure improves the epitaxial orientation of Cr2S3-2D and gradually transforms it into thicker islands. The onset of this process is visualized in Figure 2(c) and (d), obtained after annealing to 950 K. The LEED pattern of Figure 2(c) displays now a pronounced modulation of the ring intensity. Spot-like ring maxima are aligned to the first-order Gr reflections while the ring intensity in between has faded away. The STM topograph in Figure 2(d) displays now extended Cr2S3-2D islands with less wavy steps. One island, marked by a green arrow, has a somewhat different appearance and makes straight edges with angles of 60∘ or 120∘ at its corners. As shown by the profile along the white line in Figure 2(d), this island displays a uniform apparent height of nm at V, i.e., its apparent height is about 0.23 nm larger than Cr2S3-2D islands. The atomically resolved inset of the higher island makes plain that the structure has also changed. Compared to Cr2S3-2D a trimerization of the surface atoms took place, leading to triangular depressions with the -R30∘ periodicity. The underlying lattice parameter did not change within the limits of error, i.e., . As argued below, we interpret the CrxSy-2D compound represented by the 1.02 nm high island with the superstructure in Figure 1(d) as CrS4-2D of which the side view is shown in Figure 1(e).
While phase pure Cr2S3-2D is formed up to 800 K, upon increasing the annealing temperature, the transformation of Cr2S3-2D is not limited to CrS4-2D, but thicker and thicker layers are formed [see Figure S1 in the Supporting Information (SI)]. In the absence of an S vapor during annealing or for larger initial deposited amounts, the transformation of Cr2S3-2D to thicker layers sets in already at lower temperatures (see Figure S2 and Figure S3 of the SI). While the CrxSy-2D materials thicker than CrS4-2D are beyond the scope of the present manuscript, we point out that the apparent heights for V increase in the sequence 0.79 nm, 1.02 nm, 1.29 nm, 1.56 nm, 1.82 nm and so forth (see Figure S3), i.e., the levels above Cr2S3-2D increase in steps of 0.23 nm to 0.27 nm.
A key piece of information still missing in order to pinpoint the structure of Cr2S3-2D is its composition. To this end, we conducted a set of Cr calibration experiments, of which one is visualized by Figure 3. After 210 s of Cr deposition the total coverage of pseudomorphic single and double layer Cr islands on Ir(111) was 0.39 monolayers as shown in Figure 3(a). That is 0.39 of the surface atomic density of Ir(111) being atoms/m-2. Next, Cr2S3-2D was formed by Cr deposition with the same evaporator for exactly the same time and with exactly the same settings in a S background pressure on Gr/Ir(110) and subsequent annealing to 750 K. As shown in Figure 3(b) the resulting coverage of Cr2S3-2D is 0.31 of the surface area. Considering the different lattice parameters of Ir(111) [0.2715 nm] and Cr2S3-2D [0.341 nm], we arrive at a Cr content of Cr atoms per Cr2S3-2D unit cell. For a single layer of CrS2 one would have expected 1 Cr atom per unit cell, not 2.
Both phases, Cr2S3-2D and CrS4-2D, can be grown phase pure with nearly complete coverage. In a growth experiment similar to the one represented by Figure 2(a) and (b), but using a larger Cr amount, a coverage of 0.85 Cr2S3-2D is achieved [see Figure 4(a)]. The layer is fully coalesced, and only some holes are left. The height profile in the lower panel of Figure 4(a) and the LEED pattern of Figure 4(b) are consistent with our assignment. The Cr deposition time was 360 s.
After preparation of the Cr2S3-2D represented by Figure 4(a) and (b) we conducted a second growth step involving additional Cr deposition for 150 s with identical evaporation parameters. The resulting sample is represented by Figure 4(c) and (d). The coverage increased to 0.92 of the surface area, and the height profile in the lower panel of Figure 4(c) displays uniformly the characteristic height of CrS4-2D around 1.02 nm. The presence of CrS4-2D is corroborated by a second LEED-ring with a radius corresponding to a -R30∘ superstructure and highlighted in Figure 4(d) by a blue arrow.
Considering that Cr2S3-2D contains 2 Cr atoms per unit cell as established above, by taking into account the deposition times and coverages measured, we obtain a Cr content of for CrS4-2D. The experimentally found Cr content is thus consistent with CrS4-2D containing or 2.67 Cr atoms per unit cell.
Based on the experimental information and considering that CrxSy-2D phases are alternating sequences of S and partial or complete Cr planes 43, we are ready to make an informed hypothesis about the structure of the two CrxSy-2D phases. For Cr2S3-2D we assume a structure as depicted in Figure 1(d). It contains 2 Cr atoms per unit cell, as found experimentally. In principle, one S plane sandwiched between two Cr planes would be consistent with our experiments. However, the reactivity of the metal and growth under sulfur-rich conditions excludes the presence of bare metal at the Gr interface (see ref. 55) or the vacuum. Therefore, we assume the presence of two additional sulfur planes, making the compound inert and resulting in the stoichiometry Cr2S3-2D. Although we cannot guess the precise stacking, it is plausible to be similar to the bulk compounds Cr2S3 or Cr5S6 [Figure 1(a) and (c)], when disregarding the presence of Cr vacancies in these compounds. The observed hexagonal surface symmetry with a lattice parameter of 0.341 nm matches reasonably with the lattice parameter of the Cr deficient bulk compounds Cr2S3 and Cr5S6, which is in the range from 0.343-0.345 nm. Previous DFT calculations for Cr2S3-2D yield lattice parameters of 0.343-0.348 nm 39, 40 in acceptable agreement with our experiment. We emphasize that the absence of any superstructure together with 2 Cr atoms per unit cell is strong evidence for complete Cr planes in Cr2S3-2D.
An additional partial or complete Cr plane plus a saturating S plane has for bulk CrxSy materials an average height of 0.28 nm 43, in close agreement with the step height of 0.26-0.27 nm measured for thicker CrxSy-2D by us. The slightly smaller apparent step height between Cr2S3-2D and CrS4-2D of 0.23 nm fits reasonably well, and the slight deviation may be of electronic origin. It is, therefore, most plausible to assume now the presence of an additional 2/3 Cr plane (based on the measured additional Cr amount of 0.62 of a full plane) and an additional S plane giving rise to the proposed stoichiometry CrS4-2D. The presence of a 2/3 Cr plane within the stack provides a natural explanation for the observed -R30∘ superstructure, although the position of the 2/3 Cr plane is uncertain. Based on the surface sensitivity of LEED signals, it is unlikely that the 2/3 Cr plane is the lowermost in the stack. Whether the partial Cr plane is in the center as indicated in Figure 1(e) or at the top cannot be decided based on the experimental data provided up to now.
Lastly, the stability and inertness of the covalently bound Cr2S3-2D is highlighted through Figure 5. As documented by the two subsequent STM topographs of Figure 5(a) and (b), a large Cr2S3-2D island can be laterally shifted by the STM-tip interacting with the island edge without significant defect formation. This is a direct proof of just a weak van der Waals interaction between Cr2S3-2D and Gr/Ir(111). Figure 5(c) displays an STM topograph of Cr2S3-2D islands taken in a vacuum after exposure for three hours to ambient conditions [see Figure S6 for a large-scale topography of the same sample]. Not unexpectedly, adsorption on Gr and Cr2S3-2D is apparent, but the structural integrity of the islands is obvious.
2.2 Absence of CrS2
Despite substantial efforts, the growth of CrS2 using our molecular beam epitaxy approach remained unachievable, both as mono- as well as multilayer. In all experiments, including those with variations in growth and annealing procedure, we always observed a plain hexagonal surface lattice with 0.341 nm as the lattice parameter. This rules out the formation of the predicted orthorhombic CrS2 phases 33, 52, the formation of the T’-phase 25, 31, and distorted T-phases 36, 31. H- and undistorted T-CrS2 display a hexagonal lattice, but with a significantly smaller lattice parameter of nm 31, 25, 34, 36.
Moreover, for the H-,T-, and T’-phases of CrS2 one would expect a significantly smaller single layer height than the 0.79 nm measured for Cr2S3 at our reference voltage. This is because CrS2 is a three-plane S-Cr-S system rather than a five-plane S-Cr-S-Cr-S system as for Cr2S3-2D. Based on the height difference between Cr2S3-2D and CrS4-2D of 0.23 nm, one would expect a single layer of CrS2 to display an apparent height well below 0.60 nm. However, this was never the case and at the reference voltage we invariably measured an apparent height of 0.79 nm.
Growth variations involved lowering [see Figure S5(a)] and increasing the sulfur pressure during growth and annealing, as well as changing the growth temperature in the range from 160 K [see Figure S5(b)] to 600 K [see Figure S5(c)]. Particularly, we also changed from Gr/Ir(110) to Gr/Ir(111) as growth substrate [see Figure S6], without being able to detect an indication for CrS2 formation. Noteworthy, when grown on Gr/Ir(111) the apparent heights of the islands at V are 0.86 nm and 1.09 nm for Cr2S3-2D and CrS4-2D, respectively. This increase by about 0.07 nm compared to Gr/Ir(110) is primarily attributed to a larger van der Waals gap between CrxSy-2D and Gr. In fact, Gr is physisorbed on Ir(111) with weak chemical modulation 56, whereas it is chemisorbed on Ir(110) 54. This leads to stronger interactions with CrxSy-2D in the latter case and consequently to a reduced height above Gr.
The MBE growth method used here has so far enabled us to reproduce whatever transition metal disulfide described in the literature that it was applied to 57, 58, 55. Therefore, the inability to grow CrS2-2D using the same method indicates that it may be worthwhile to reevaluate the evidence for the existence of CrS2.
TEM images of CrS2 are provided in four publications 25, 26, 27, 28 and display a perfectly hexagonal lattice ruling out the observation of orthorhombic, distorted T- or T’-phases. Based on the specified net plane distances or scale bars provided, the lattice parameter is in the range of 0.33-0.36 nm, in descent agreement with the lattice parameter of 0.341 nm measured for Cr2S3-2D and CrS3-2D, but at variance with the calculated lattice parameters of nm 31, 25, 34, 36, 37 for the undistorted T- or H-phase. The measured CrS2 lattice parameters also agree with the reported lattice parameters of Cr2S3 thin films [see Figure 1(c)] ranging from 3.4 Å to 3.51 Å18, 22, 16, 17.
The minimum heights of CrS2 flakes measured by atomic force microscopy are 2 nm 26, 1.5 nm 27, 0.75 nm 25, and 0.682 nm 28. Despite the three significant digits provided in the latter height, the noise level on each terrace is 0.2 nm. Given the experimental uncertainties, the measured heights are also compatible with Cr2S3-2D, or CrS3-2D and Cr2S3 thin films for the thicker layers.
The characteristic Raman peaks of CrS2 scatter with a few wave numbers around 175 cm-1, 252 cm-1, 283 cm-1, and 359 cm-1 26, 27, 25, 28. It also has to be noted that Raman peaks are measured for thicker CrS2 films for signal reasons 27. Peak number and positions agree well with the ones reported for Cr2S3 thin films, specifically when considering their thickness-dependent variability 18, 19, 17.
XPS characterization through the core level doublets S 2p3/2/S 2p1/2 and Cr 2p3/2/Cr 2p1/2 does neither provide a clear distinction between between CrS2 25, 26, 28and thin films of Cr2S3 18, 16, 20, 22, 17, 21. For instance, the data provided by Nair et al. 26 for CrS2 agree perfectly with the one provided by Cui et al. 21 for Cr2S3 thin films. Although XPS can be used for composition analysis, precise composition analysis is not feasible without proper calibration, taking depth-dependent attenuation of the signal into account and considering the depth distribution of the species in a layered material. It also has to be noted that the top three layers of S-Cr-S in thin Cr2S3 films and CrS2 are identical. Lastly, the prominent intensity peak at in X-ray diffraction of CrS2 25, 28, 27 is also present for thin Cr2S3 films 16, 20.
Given the absence of CrS2 as a bulk material and the ambiguity surrounding the evidence for its existence as a 2D material, it may be valuable to consider how to obtain more conclusive evidence of CrS2.
2.3 Electronic structure from experiment
The electronic structure of Cr2S3-2D and CrS4-2D was investigated by STS. Figure 6(a) and 6(b) present differential conductance () spectra of Cr2S3-2D and CrS4-2D in the energy range . The spectra were measured far away ( nm) from island edges and are highly reproducible. The most remarkable feature in both spectra is the band gap, which seems to range from the Fermi level at = 0 V to the valence band edges (VBEs) at V and in Cr2S3-2D and CrS4-2D, respectively. While high-resolution spectra confirm the sharp rises at negative voltages and thus the assignment of the VBEs, the apparent conduction band edges (CBEs) at the Fermi energy are suspicious. High-resolution spectra near shown as insets in Figure 6(a) and 6(b) make plain that there is a finite density of states below .
Such a smeared-out behavior of the conduction band edge (CBE) has been observed in metallic MoS2 (see Figure S5 of the Supplementary Data of ref. 59). Detection of a sharp CBE for MoS2 in spectra was hampered by the fact that the conduction band minimum is located at the K-point of the first Brillouin zone (BZ). For such electrons with a large parallel momentum STS is relatively insensitive. As a result, precise energy assignment to the CBE is challenging. Here we assume a similar origin of the smeared-out CBE, which would imply a large parallel momentum of the electrons at the conduction band minimum. As an estimate, we suggest the CBE at V for both materials, yielding approximately band gaps of eV and eV for Cr2S3-2D and CrS4-2D, respectively. Furthermore, both spectra exhibit a pronounced dip at . Features at , including dips and peaks, are commonly observed in spectra of 2D materials 60 and are likely related to electronic correlations. However, the exact origin remains uncertain in this instance.
Whether a 2D material with a full band gap is a semiconductor or a metal depends on the position of which may be affected by the substrate. In MoS2, a semiconductor-metal transition was driven by doping-induced work function lowering of the Gr substrate 59. Vacuum level alignment of Gr and MoS2 then implied a downshift of the MoS2 band structure until the CBE moved into the Fermi level and an interface dipole was formed by charge transfer from Gr to MoS2. We consider here for Cr2S3-2D and CrS4-2D a similar scenario. That is while the freestanding materials might be semiconducting, we speculate that a lower Gr work function implies, through vacuum level alignment for the CrxSy-2D compounds, could lead to a downward shift of their bands and, eventually, the formation of an interface dipole due to charge transfer into their conduction bands. Consequently, they become metallic. Such a scenario would imply that the work function of the CrxSy-2D compounds to be larger than the one of Gr. To substantiate this speculation, we estimated the work functions of the CrxSy-2D compounds using the method of field emission resonances 61, 62 and found them in fact to be larger by more than one electronvolt compared the one of Gr/Ir(110) (see Figure S7 of the SI).
A notable feature in the spectrum of and CrS4-2D is a sharp rise peak at the VBE as highlighted by the inset in the lower left corner of Figure 6(b). Such a sharp feature might be related to a van Hove singularity of a surface state 63.
Work function differences and the density of states also affect the measured apparent height in STM. The apparent heights of Cr2S3-2D and CrS4-2D were up to now only specified for V. This choice is meaningful because at positive bias the apparent heights are stable and vary only marginally with , as expected for a metal, for which the apparent height usually agrees rather well with the geometric height. Not unexpectedly, at negative bias in the range of the band gap, apparent heights up to 0.15 nm lower are measured (see Figure S8 in the SI).
Finally, it is found that the electronic structure remains invariant with respect to the substrate, whether Gr/Ir(110) or Gr/Ir(111), as evidenced by comparison in Figure Figure S9. The invariance of the measured electronic structure is consistent with a weak interaction and the absence of significant hybridization with the two different substrates.
2.4 Density functional theory calculations
Based on the experiments, a structure S-Cr-S-Cr-S consisting of five atomic planes is inferred for Cr2S3-2D, but the stacking sequence could not be established. In our DFT calculations, we tested the typical trigonal prismatic (1H) and octahedral (1T) coordination of Cr with respect to S, with stackings displaying aligned and shifted Cr positions between the two Cr planes (see Table S1 of the SI). To identify the most stable structure for Cr2S3-2D among the resulting eight possible crystal configurations, their total energies were calculated by DFT. The results, presented in Table S1 of the SI, give the lowest energy for the crystal structure with aligned 1T/1T Cr coordination, as illustrated in Figure 7(a) to (c). This crystal structure corresponds to the classical NiAs-structure and is the same as for bulk Cr2S3 and Cr5S6, when disregarding the presence of Cr vacancies in the bulk structures. The NiAs-structure of Cr2S3-2D is preferred by 0.17 eV per formula unit compared to a structure with the same octahedral coordination of the Cr atoms, but shifted Cr positions on the atomic planes, and by more than 0.6 eV compared to all other structures. The calculated lattice parameter of 0.344 nm agrees reasonably well with the experimental value of 0.341 nm. Furthermore, no superstructure is observed, consistent with the experimental results.
The calculated DFT (for Hubbard ) band structure is shown in Figure 7(d). It exhibits an indirect band gap of 0.9 eV in reasonable agreement with our experimental estimate of 1.0 eV. The conduction band minimum is located between the K and M points of the first BZ. The large parallel momentum of the electrons at the conduction band minimum is consistent with the insensitivity of STS to its position and its smeared-out appearance in the spectra [see Figure 6(a)]. The valence band maximum lies close to the center () of the first BZ. Correspondingly, the smaller parallel momentum makes plain why the VBE appears rather sharp in the Cr2S3-2D spectrum of Figure 6(a).
Recalculating the band structure with does not change the overall picture, but the magnitude of the band gap is affected. It increases with increasing , reaches a maximum around 1.32 eV for eV and then drops to 0.84 eV for eV. Taking into account spin orbit coupling (SOC) decreases the bandgap size only slightly by 20 meV, for the complete list of bandgap sizes as a function of see Table S2 in the SI.
A metallic state of Cr2S3-2D is obtained when the calculations are carried out considering the Gr substrate. Indeed, the CBE shifts below , indicating a semiconductor-metal transition of Cr2S3-2D (see projected density of states (PDOS) in Figure S10 in the SI).
Next, we assess the magnetic properties of the Cr2S3-2D structure via DFT calculations. We find that the magnetic moment of the Cr+3 atoms (i.e., 3 ) is affected by SOC and reduced to 1.73 per Cr atom, indicating the presence of magnetism in the structure. To infer the magnetic ground state of Cr2S3-2D, intralayer Cr-Cr coupling was considered ferromagnetic (FM). This assumption is consistent with expectation based on the Goodenough-Kanamori-Anderson rule, which favors FM coupling when the bond angle Cr-S-Cr is close to 90° (in Cr2S3-2D it is 85.77°). For the interlayer coupling, we considered FM or A-type (or layerwise) antiferromagnetic (AFM) states, for a schematic representation, see Figure S11 in the SI. The energy difference between these two states is meV, that is the A-type AFM configuration is energetically preferred.
The interlayer magnetic exchange splitting calculated by GGA is notably weak, emphasizing the importance of accurately accounting for electronic correlation effects when estimating magnetism. To address this, we analyzed the energetics of FM and AFM states by varying the parameters in GGA+U calculations. As the on-site Coulomb term increased, a transition from the AFM to the FM state was observed, as shown in Table S2 in the SI. This shift occurs because the super-exchange interlayer coupling is inversely related to the hopping parameter; thus, raising reduces the stability of AFM coupling relative to FM coupling.
For CrS4-2D not only the stacking sequence, but also the position of the fractional Cr plane was uncertain from experimental data. Experimentally, LEED ruled out the fractional Cr plane to be the lowermost. Since for Cr2S3-2D octahedral (T) coordination of Cr atoms was preferred, we computed for CrS4-2D the four different structures consistent with these two constraints as outlined in Table S3 of the SI. We find the minimum energy configuration to be a NiAs-type structure with the fractional Cr plane in the center, as represented in Figures 8(a) to (c). This structure is preferred by 0.6 eV per formula unit. The DFT optimized configuration reflects the Cr5S6 motif of intermediate Cr layers filled only by [see Figures 1(c) and 1(e)]. The relaxed structure has a lattice constant of 0.347 nm, which agrees with an acceptable error to the experimental 0.341 nm. The thickness difference between the relaxed crystal structures of Cr2S3-2D and CrS4-2D is 0.26 nm, a value that closely aligns with the apparent height difference of 0.23 nm observed in STM.
The ()-R30° superstructure visible in LEED and STM is according to our DFT calculations due to distortions caused by the missing Cr atoms in the middle Cr plane. The missing Cr atoms affect the position of S in the topmost plane. For three S atoms surrounding one Cr vacancy their displacements are highlighter by arrows in Figure 8(a). The DFT simulated constant-current STM topograph of CrS4-2D shown in the upper panel of Figure 8(d) closely resembles the experimental STM constant-current topograph shown in the lower panel.
Band structure and PDOS of CrS4-2D are presented in Figure 8(e,f). The unfolded band structure has the conduction band minimum at the boundary of the original first BZ at the M-point, consistent with the smeared out CBE observed in the spectra in Figure 6(b) and as discussed above. The valence band maximum is located roughly in the middle between the M- and the -point. Thus, the electrons at the VBE have smaller parallel momentum, consistent with the sharp VBE seen in STS. The calculated band gap of CrS4-2D is with 0.33 eV clearly smaller than the one of Cr2S3-2D, consistent with experiment. However, a quantitative agreement of the band gap with the experiment is not expected. When is increased in the calculations, also the band gap increases and for a reasonable value of eV it amounts to 0.62 eV, close to the experimental result (see Table S4 of the SI).
For CrS4-2D, the oxidation state of Cr atoms is also thanks to the identical stoichiometry of the 2D compound. The magnetic moment is retained and takes identical values to Cr2S3-2D, i.e., 3 without and 1.73 per Cr atom with SOC. To explore the magnetic ground state of CrS4-2D, we assumed similarly to Cr2S3-2D FM intralayer coupling. For the interlayer magnetic order three configurations were considered: one FM and two ferrimagnetic (FiM) ones. Two adjacent FM coupled Cr planes (one outer and the inner) couple AFM to the third (FiM-I), or the two outer Cr planes are both coupled AFM to the inner partial Cr plane (FiM-II), (see Figure S12). The relative energies between FM and FiM-I/FiM-II configurations are listed in Table S4. The results indicate that the ground state of CrS4-2D is FiM-I. CrS4-2D has an odd number of Cr planes and is a highly uncompensated antiferromagnet, as the magnetic moment of both outer Cr planes point in the same direction, while the fractional Cr plane contains magnetic moments pointing in the opposite direction. It results in ferrimagnet behavior displaying clearly spin-polarized bands, best visualized in Figure 8(e). Interestingly, CBE is dominated by spin-up states and VBE is composed of spin-down electrons. Like Cr2S3-2D, the interlayer magnetic coupling of CrS4-2D changes from AFM to FM as the is considered. The relative energies between FM and FiM-I/FiM-II configurations as a function of are given in Table S4 of the SI.
3 Conclusion
We demonstrate the growth of two new chromium-sulfur 2D materials, Cr2S3-2D and CrS4-2D, in single unit cell thickness. Cr2S3-2D exhibits threefold symmetry with an in-plane lattice constant of nm, an apparent STM height of 0.79 nm, and 2 Cr atoms per unit cell. Upon annealing to 850 K or growing over the complete layer, Cr2S3-2D transforms into CrS4-2D with the same lattice parameter, but a thickness of 1.02 nm. This transition results in an additional -R30∘ superstructure in respect to the original lattice. Both materials can also be grown phase pure to almost a full layer. Cr2S3-2D and CrS4-2D display a NiAs-type structure terminated by S planes. However, the middle Cr layer in CrS4-2D displays Cr vacancies with 1/3 of the Cr positions missing, resulting in a ()-R30° superstructure.
Despite considerable variations in the MBE growth procedure, none of the predicted CrS2 phases could be synthesized. Based on a critical assessment of the experimental evidence for the existence of CrS2 and considering the absence of CrS2 as bulk material, it is concluded that additional evidence may be needed to make the existence of CrS2 unambiguous.
Our DFT calculations predict that both 2D materials are indirect bandgap semiconductors in their freestanding form. Interestingly, we find that when supported on Gr, they exhibit metallic behavior due to their relatively large work function compared to Gr. DFT calculations find sizable magnetic moments for both phases, possessing magnetic order with coupling depending on the degree of Coulomb repulsion, a characteristic that requires experimental validation. The phase-pure, large-area growth of Cr2S3-2D and CrS4-2D enables detailed investigation of their magnetic properties using averaging techniques and establishes these compounds as promising 2D magnetic materials for potential applications.
4 Experimental Section
4.0.1 Experimental methods
In situ growth and analysis of CrxSy-2D were performed in a variable temperature and a low temperature bath cryostat ultra-high vacuum STM system, both with base pressure of mbar. Each system is equipped with standard sample preparation tools, MBE growth facility, LEED or microchannel plate LEED, and STM operating down at 300 K and at 1.7 K.
4.0.2 Substrate preparation
Ir(110) is cleaned and prepared in its unreconstructed state by cycles of 4.5 keV Xe+ ion sputtering and flashing annealing to 1510 K followed by cooling to 400 K in mbar oxygen pressure. Ir(111) is cleaned by cycles of 1 keV Ar+ ion sputtering and flash annealing to 1510 K.
To grow a single crystal Gr on Ir(110), clean unreconstructed Ir(110) is heated to 1510 K and exposed at this temperature for 240 s to mbar ethylene 54. To grow single crystal Gr on Ir(111), cleaned Ir(111) is exposed to mbar ethylene at room temperature for 120 sec. After flash annealing to 1470 K without ethylene, subsequently the sample is exposed to mbar ethylene at 1370 K for 600 sec for completion of Gr growth 64. The quality of the as-grown Gr on Ir(110) and Ir(111) is checked by LEED and STM.
4.0.3 Sample preparation
To grow CrxSy-2D, Gr/Ir(110) is exposed to a Cr flux in the range of atoms m-2 s-1 from an e-beam evaporator simultaneously with a sulfur flux obtained by decomposing pyrite (FeS2) in a Knudsen cell in a distance of 8 cm from the sample. The S flux is characterized by a distant ion gauge measuring a S pressure of mbar. This growth step was performed typically at 300 K, if not specified differently. To complete CrxSy-2D formation, the sample is subsequently annealed for half the initial deposition time in the same sulfur flux at temperatures ranging from 500 K to 1050 K and specified in the respective captions.
4.0.4 LEED measurements
LEED patterns were acquired at 300 K. Distortions in the reciprocal space due to the flat microchannel plate are corrected for the microchannel plate LEED using a Phython script.
4.0.5 STM measurements
STM data is acquired at 300 K in the variable temperature STM system and at 1.7 K in the bath cyrostat STM system. Constant current topographies are presented with sample bias and tunneling current specified in the captions. Image processing was conducted using WSxM 65.
4.0.6 STS measurements
For STS, W tips were further prepared on Cu(111) until the signature of the Cu(111) surface state was well reproduced. The spectra are recorded at 1.7 K with stabilization bias and stabilization current using a standard lock-in technique with a modulation frequency and modulation voltage with respective values specified in the captions.
4.0.7 Theoretical calculations
The energetics and magnetic properties of all considered systems were investigated using spin-polarized density functional theory (DFT) as implemented in the VASP code 66, 67. All the calculations were carried out using the Perdew-Burke-Ernzerhof exchange-correlation functional 68. To account for the high correlation effects among d-orbital electrons of chromium atoms, an effective Hubbard value () was considered for Fe atoms within the DFT + U approach. A plane-wave cut-off energy of 500 eV and force tolerance of 0.01 eV/Å was set for geometry optimization. Van der Waals interactions were considered using the DFT-D2 method proposed by Grimme 69. A vacuum space of approximately 30 Å was introduced to model the slabs between the periodic images in the confinement direction.
The magnetic configuration does not influence the structure outcome, as the magnetic order energy dependence is an order of magnitude smaller than the stacking dependence. The magnetic order was calculated only for the NiAs-type structure.
Note that, due to the minimal contribution of the SOC effect to the total energy, structural changes arising from the SOC effect were not included in the calculations. Furthermore, the energy differences among magnetic states typically change by less than 11 meV with the SOC effect. Therefore, SOC effects are unlikely to significantly alter the results presented in this work.
Funding from Deutsche Forschungsgemeinschaft (DFG) through CRC 1238 (project number 277146847, subprojects A01 and B06) is acknowledged. J.F. and W.J. acknowledge financial support from the DFG through project FI 2624/1-1 (project No. 462692705) within the SPP 2137 and project JO 1972/2-1 (project No. 535290457) within SPP 2244, respectively. A.V.K. acknowledges funding from DFG projects KR 4866/11-1, and the collaborative research center “Chemistry of Synthetic 2D Materials” CRC-1415-417590517. Generous CPU time grants from the Technical University of Dresden computing cluster (TAURUS) and Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu), Supercomputer HAWK at Höchstleistungsrechenzentrum Stuttgart (www.hlrs.de), are greatly appreciated.
Conflict of Interest
The authors declare no conflict of interest.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.